OpenStudy (anonymous):

163.35 is 45% of what number> How do I solve w/o a calculator?

OpenStudy (anonymous):

363

OpenStudy (anonymous):

163.45 = 0.45 x number Required Number = 163.45/0.45

OpenStudy (anonymous):

cross muliply= 163.35/x=45/100

OpenStudy (anonymous):

both mean the same wilverine

OpenStudy (anonymous):

Know a method of solving w/o a calc? I could just multiply, but another method would help

OpenStudy (anonymous):

Wait I can cross multiply this?

OpenStudy (anonymous):

that is without using calculator.. simple long division and multiplcation

OpenStudy (anonymous):

yeah the fractions is equal to each other because 163.35/x where x is 100% and 163.35 is 45% of x

OpenStudy (anonymous):

Can I cross multiply this also? 105=y*24%?

OpenStudy (anonymous):

yeah. 105/y=24/100.. but it can also solve for y, because to change a percent to a number just move decimal over 2 places to the left, 105=y*.24